# Monte Carlo Simulation Development and Implementation of the GiBUU Model for Neutrino Experiments

### v2

@@ V Results and Discussion
By investigating different particle energy thresholds and particle multiplicities, we observed that GENIE produces a higher proton multiplicity than GiBUU. For example, as shown in Figure 6, GENIE has a lower composition of events with no protons below the energy threshold in the $\nu_{\mu}CC0\pi^{\pm/0}$ topology. We also observe different neutrino interaction compositions in each model, particularly in MEC and quasi-elastic production. These differences, along with variations in nuclear medium modeling, may explain the discrepancy in the $\nu_{\mu}CC0P0\pi^{\pm/0}$ topology between the two models, which exhibited the largest difference in the double ratio from Figure 5.
@@ VI Conclusion
In conclusion, implementing the GiBUU model as a Monte Carlo has yielded valuable insights into the complex dynamics of neutrino scattering processes. This Monte Carlo reproduces correctly the GiBUU model, and it is the first step in incorporating uncertainties to be propagated into experimental data analysis. Notably, a comparative study with GENIE reveals nuanced differences, emphasizing the necessity of employing multiple models to interpret neutrino scattering data comprehensively. The significant differences observed between GiBUU and GENIE, particularly in expected topological event samples and proton multiplicity (Figure 56, and  7), highlight the importance of understanding these models to better assess uncertainties when reconstructing neutrino energy, especially for oscillation experiments. GiBUU offers a different approach to other generators with a unified nucleus description, bringing new opportunities to better explore our data and develop model uncertainties. The same methodology we used to incorporate GiBUU as an unweighted event generator could be used for other cross-section models not yet available as event generators, such as some Beyond the Standard Model theory simulations. Our simulation study employed the described technique to simulate GiBUU events within the LArBox geometry, utilizing the BNB flux.

### v3

@@ IV Ongoing Developments in Systematic Uncertainty Estimation for GiBUU
Systematic uncertainties arising from non-resonant pion backgrounds and other reaction processes, such as two-particle two-hole interactions and those specific to hadronization modeling, are not explicitly addressed here but will need to be developed to assess GiBUU’s systematic effects. Conversely, processes modeled by GiBUU, like the non-resonant pion background, could provide valuable insights into uncertainties resulting from the absence of these interactions in other neutrino event generators, such as GENIE, particularly in the context of DUNE.
@@ V Results and Discussion
By investigating different particle energy thresholds and particle multiplicities, we observed that GENIE produces a higher proton multiplicity than GiBUU. For example, as shown in Figure 6, GENIE has a lower fraction of events with no protons below the energy threshold in the $\nu_{\mu}CC0\pi^{\pm/0}$ topology. Additionally, significant differences are observed in the descriptions of the models of neutron multiplicities, as illustrated in Figure 7, with large discrepancies between the generators. We also note distinct neutrino interaction compositions in each model, particularly in MEC and quasi-elastic production.
One potential cause of these discrepancies is that GENIE does not include Pauli blocking in its treatment of FSI, resulting in many protons that should be absorbed in the nucleus being incorrectly retained. Furthermore, the treatment of FSI differs substantially between the two generators. These differences, along with other variations in nuclear medium modeling, likely explain the discrepancy in the $\nu_{\mu}CC0P0\pi^{\pm/0}$ topology between the models, which exhibited the largest difference in the double ratio shown in Figure 5.
[…]
@@ VI Conclusion
In conclusion, implementing the GiBUU model as a Monte Carlo has yielded valuable insights into the complex dynamics of neutrino scattering processes. This Monte Carlo reproduces correctly the GiBUU model, and it is the first step in incorporating uncertainties to be propagated into experimental data analysis. Notably, a comparative study with GENIE reveals nuanced differences, emphasizing the necessity of employing multiple models to interpret neutrino scattering data comprehensively. The significant differences observed between GiBUU and GENIE, particularly in expected topological event samples and proton multiplicity (Figure 56, and  8), highlight the importance of understanding these models to better assess uncertainties when reconstructing neutrino energy, especially for oscillation experiments. GiBUU offers a different approach to other generators with a unified nucleus description, bringing new opportunities to better explore our data and develop model uncertainties. The same methodology we used to incorporate GiBUU as an unweighted event generator could be used for other cross-section models not yet available as event generators, such as some Beyond the Standard Model theory simulations. Our simulation study employed the described technique to simulate GiBUU events within the LArBox geometry, utilizing the BNB flux.